OpenStudy (anonymous):

Which set of steps can you use to solve three x minus two equals thirteen in two steps?

OpenStudy (dan815):

3x-2=13

OpenStudy (anonymous):

3x-2=13 1. Add 2 to each side 3x=15 2. Divide each side by 3 x= 15/3 = 5

OpenStudy (dan815):

add 2 to both sides and divide by 3